OpenStudy (pulsified333):

Given that: f(−4)=4, f′(−4)=3, g(−4)=8, g′(−4)=9, calculate the following: (fg)′(−4)= (f/g)′(−4)= (g/f)′(−4)=

OpenStudy (pulsified333):

@mathmale

zepdrix (zepdrix):

By product rule we have,\[\large\rm (fg)'=f'g+fg'\]Evaluated at x=-4 gives us,\[\large\rm (fg)'(-4)=f'(-4)g(-4)+f(-4)g'(-4)\]Ya? :) Any confusion there?

OpenStudy (pulsified333):

a little

OpenStudy (pulsified333):

oh wait

OpenStudy (pulsified333):

(3*8)+(4*9)

OpenStudy (pulsified333):

is that right?

zepdrix (zepdrix):

Good good good.

zepdrix (zepdrix):

For the others, remember your quotient rule?

OpenStudy (pulsified333):

a little bit

OpenStudy (pulsified333):

Medals 1 ((3*8)-(4*9))/(8^2)

OpenStudy (pulsified333):

?

zepdrix (zepdrix):

\[\large\rm \left(\frac{f}{g}\right)'=\frac{f'g-fg'}{g^2}\] Mmm yes looks good!

OpenStudy (pulsified333):

how would you do the last one thought, im a little confused?

zepdrix (zepdrix):

Umm you just follow your rule for quotient.\[\large\rm \left(\frac{g}{f}\right)'=\frac{g'f-gf'}{f^2}\]Dtop*bottom minus Dbottom*top all divided by bottom^2